OpenStudy (timaashorty):

If y is a real number and 4-y/3+y is undefined, what is the value of 3+y/4-y? A. 6 B. 4 C. 3/4 D. 0 E. undefined

OpenStudy (nikato):

To make the first equation undefined, the denominator must equal 0. So 3+y=0 -3. -3 ---------- y=-3 Plug this into the second equation 3+(-3) /4-(-3) =0/7 =0

OpenStudy (timaashorty):

Thank you(:

OpenStudy (nikato):

No problem